Convert 80mL (4.03g) of acetic acid into moles of acetic acid


1
Expert's answer
2022-06-06T08:25:03-0400

n(CH₃COOH) = m(CH₃COOH) / Mr(CH₃COOH)

n(CH₃COOH) = 4.03 g / 60,05 g/mol = 0.067 moles


Cm = 0.067 moles * 0.08 L = 5.36 * 10-3 M